OpenStudy (anonymous):

Get a medal. Simple.....question.......I think...... Solve 4/3 = 12/x A. 9 B. 48/3 C. 8/3 D. 11

OpenStudy (the_fizicx99):

I think A

OpenStudy (the_fizicx99):

To isolate x transform 4/3 = 12/x to 12/x = 4/3 Do the reciprocal of both sides: x/12 = 3/4 Then just solve for x (multiply both sides times 12)

jigglypuff314 (jigglypuff314):

4/3 = 12/x multiply both sides by x 4/3 x = 12 then multiply both sides by 3 4x = 12 * 3 then divide both sides by 4 so x = (12 * 3) / 4 =

OpenStudy (anonymous):

Thanks so much *hug*

OpenStudy (the_fizicx99):

-hugs back- :* muaa xD